OpenStudy (anonymous):

It looks right to me, so I must be wrong... help? :') 1. Find, explain, and correct the error in the solution given. (1) (2) (3) (4) (5) (6) (7) 3x = 24 (8) (9) x = 8

OpenStudy (blurbendy):

why are 1-6, and 8 empty?

OpenStudy (anonymous):

Whoops... for some reason half of the thing is missing O_O they do have weird symbols, I guess my computer isn't capable of loading them.... brb, let me type it manually

OpenStudy (blurbendy):

you can also take a screenshot and post it here using Attach File

TheSmartOne (thesmartone):

or using lightshot...

OpenStudy (anonymous):

Ohhh, good idea! thanks! Here you go :)

OpenStudy (anonymous):

It's combined operations, btw, if that helps c:

OpenStudy (blurbendy):

Notice anything funny about step 4?

OpenStudy (blurbendy):

remember PEMDAS Parentheses Exponents Multiplication Division Addition Subtraction as the order of operations

OpenStudy (mertsj):

\[\frac{3x-2}{4}=\frac{3x}{4}-\frac{2}{4}=\frac{3x}{4}-\frac{1}{2}\]

OpenStudy (mertsj):

So, in step 4, you should add 1/2 to both sides NOT 2

OpenStudy (anonymous):

Okay, thank you! :D

OpenStudy (blurbendy):

if you dont want to deal with fractions, you can also just multiply both sides by 4